OpenStudy (anonymous):

Find all of the complex cube roots 1 + i

OpenStudy (anonymous):

would you sue the nth roots theorem for this problem? or De Moivre's Theorem?

OpenStudy (anonymous):

*would you start

OpenStudy (anonymous):

first write this in polar form, that will help do you know how to find the absolute value and the angle?

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\] and in your case it is \(|1+i|=\sqrt{1^2+1^2}=\sqrt2\)

OpenStudy (anonymous):

then you need the angle

OpenStudy (anonymous):



OpenStudy (anonymous):

yes for polar i had: z= √2 (cosπ/4 + i sinπ/4)

OpenStudy (anonymous):

oooh then you are completely done almost that is the hard part

OpenStudy (anonymous):

would i use the nth roots theorem for this one? or the de moivre's theorem? i think its nth roots??

OpenStudy (anonymous):

divide the angle by 3, and take the real cubed root of the absolute value

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

nth roots?

OpenStudy (anonymous):

\[\sqrt[6]{2}\left(\cos(\frac{\pi}{12})+i\sin(\frac{\pi}{12})\right)\] for the first one

OpenStudy (anonymous):

there will be 3 answers right?

OpenStudy (anonymous):

i dunno i only ever heard of demoive, not the "nth root theorem" yeas, there are three

OpenStudy (anonymous):

once you have one you have them all, divide the circle up in to three equal parts either that, or rewrite \[\sqrt2\left(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4})\right)\] and \[\sqrt2\left(\cos(\frac{9\pi}{4})+i\sin(\frac{9\pi}{4})\right)\] and start again

OpenStudy (anonymous):

i.e add \(2\pi\) to the angle, then divide by 3 then repeat one more time to find the third one

OpenStudy (anonymous):

can i tell you wht i got for my 3 answers? and then can you tell me if theyre right?

OpenStudy (anonymous):

z1= 3.066+ i 0.822 z2= 1.587 + i 2.750

OpenStudy (anonymous):

z3= -0.822 - i 3.067

OpenStudy (anonymous):

ARE my answers right? anyone??